OpenStudy (anonymous):

Write the standard form of the equation of the conic section with the given charateristics. 1. Parabola with vertex at (0, 0) and directrix at x = 5. 2. Circle with center at (3, -1) and radius 4. 3. Ellipse with foci at (4, 2) and (10, 2) and co-vertices at (7, 0) and (7, 4) 4. Hyperbola with vertex at (0, 4), focus at (0, 5), and center at (0, 1)

OpenStudy (dumbcow):

1. x = -1/20y^2 2. (x-3)^2 + (y+1)^2 = 16 3. (x-7)^2/13 + (y-2)^2/4 = 1 4. (y-1)^2/9 -x^2/7 = 1